For the diprotic weak acid H2A, Ka1 = 3.8 × 10-6 and Ka2 = 5.8 × 10-9. What is the pH of a 0.0700 M solution of H2A? What are the equilibrium concentrations of H2A and A2- in this solution?
Answer #1

H2A = 0.0700M

Ka1= 3.8x10^-6

H2A -------------- HA- + H+

0.0700                0        0

-x                      +x         +x

0.0700-x            +x       +x

Ka1= [HA-][H+]/[H2A]

3.8x10^-6 = x*x/(0.0700-x)

for solving the equation

x= 0.00051

[H+] = 0.00051M

-log[H+]= -log(0.00051)

PH= 3.29

[H2A] = 0.0700 - 0.00051=0.0695M

[HA-] = 0.00051M

             HA- --------------- H+ + A-2

     0.00051                    0.00051

Ka2= [H+][A-2]/[HA-]

5.8x10^-9= 0.00051x[A-2]/0.00051

[A-2]= 5.8x10^-9M